### v1

# Machine Learning Nonadiabatic Dynamics: Eliminating Phase Freedom of Nonadiabatic Couplings with the State-Interaction State-Averaged Spin-Restricted Ensemble-Referenced Kohn–Sham Approach

Abstract
Excited-state molecular dynamics (ESMD) simulations near conical intersections (CIs) pose significant challenges when using machine learning potentials (MLPs). Although MLPs have gained recognition for their integration into mixed quantum-classical (MQC) methods, such as trajectory surface hopping (TSH), and their capacity to model correlated electron-nuclear dynamics efficiently, difficulties persist in managing nonadiabatic dynamics. Specifically, singularities at CIs and double-valued coupling elements result in discontinuities that disrupt the smoothness of predictive functions. Partial solutions have been provided by learning diabatic Hamiltonians with phaseless loss functions to these challenges. However, a definitive method for addressing the discontinuities caused by CIs and double-valued coupling elements has yet to be developed. Here, we introduce the phaseless coupling term, $\Delta^{2}$, derived from the square of the off-diagonal elements of the diabatic Hamiltonian in the SSR(2,2) formalism. This approach improves the stability and accuracy of the MLP model by addressing the issues arising from CI singularities and double-valued coupling functions. We apply this method to the penta-2,4-dieniminium cation (PSB3), demonstrating its effectiveness in improving MLP training for ML-based nonadiabatic dynamics. Our results show that the $\Delta^{2}$ based ML-ESMD method can reproduce ab initio ESMD simulations, underscoring its potential and efficiency for broader applications, particularly in large-scale and long-timescale ESMD simulations.
[…]
{tocentry}
1 Introduction
The study of excited state phenomena has been a significant focus for researchers across various fields, offering profound insights into processes, such as photoisomerization1, 2, 3, 4, photocatalysis5, 6, 7, solar cells8, 9, 10, and many other light-involved chemical processes11, 12, 13, 14, 15, 16. In computational chemistry, the light-matter interaction in molecular systems can be explored using excited state molecular dynamics (ESMD). ESMD provides chemical insights by allowing for time-resolved tracking of molecular systems. Unlike ground state dynamics, where the nuclear wavepacket fluctuation is confined to a single adiabatic potential energy surface (PES), ESMD considers the presence of nonadiabatic interactions across multiple PESs.
From a practical perspective, mixed quantum-classical (MQC) approaches17, 18, 19, 20, 21 are commonly used where MQC methods treat electronic propagation quantum mechanically while the nuclei move classically along multiple trajectories. Since conventional photochemical reactions occur over timescales ranging from hundreds of femtoseconds to nanoseconds, a large number of electronic structure calculations are required during an ESMD simulation. For instance, predicting PESs for 100,000 nuclear configurations is necessary to simulate dynamics for 1 ps with a time step size of 1 fs and 100 nuclear trajectories. Moreover, ESMD is highly sensitive to the accuracy of nonadiabatic coupling vectors (NACVs), particularly in cases involving conical intersections (CIs). In such scenarios, quantum chemical methods that adequately account for electronic correlation must be chosen. Although there are highly accurate quantum mechanical methods, e.g. equation-of-motion coupled cluster singles and doubles (EOM-CCSD)22, 23, MQC simulations require hundreds or even thousands of trajectories, making it computationally prohibitive to use such high-level ab initio theories, even for small systems. This makes the trade-off between accuracy and computational cost a central issue in MQC dynamics, presenting an ongoing challenge in balancing accuracy with feasibility.
Over the past decade, machine learning potentials (MLPs) have enabled us to overcome the accuracy-cost trade-off in theoretical chemistry. MLPs can predict molecular properties with an accuracy comparable to traditional ab initio theories but at a significantly lower computational cost. 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46 These advances in ML-based property prediction can be integrated into the MQC dynamics by replacing quantum mechanical calculations with MLPs. However, applying MLPs to ESMD introduces unique challenges that are not typically encountered in ground state MD simulations. A key difficulty is the prediction of NACVs near CIs of PESs. The NACV formula incorporates the energy difference between two electronic states in the denominator. Even minor energy errors can cause significant NACV inaccuracies near CIs. Additionally, NACVs are not differentiable and diverge at CIs. Various studies have proposed solutions to the singularities at CIs. The first approach is to avoid the direct prediction of NACV by using a method that does not require NACV. The simplicity of the Landau–Zener formalism47 has attracted researchers and has been employed in various studies48, 49. The Zhu-Nakamura method50, based on the Landau–Zener formalism, has also been utilized in ML-based noadiabatic dyanmics51, 52. Another approach focused on learning diabatic Hamiltonian instead of adiabatic Hamiltonian53, 54, 55. This provided smooth and differentiable predictions for energy and NACVs near CIs.
While the issue of CI singularity can be addressed through diabatic transformation, another challenge arises from the double-valued nature of the coupling elements. This double-valued nature, known as the geometric phase effect56, imparts a similar double-valued character to NACVs. As a result, NACVs derived from electronic structure calculations have arbitrary signs, leading to discontinuities that hinder efficient model training. To circumvent this problem, a phaseless loss function was introduced57 and applied in several authors of this work58. However, this approach does not fully resolve the problem, as the training set contains two values (one positive and one negative) for the same molecular geometry. A more recent work proposed an auxiliary single-valued function59, where the double-valued function, $\Gamma$, to eliminate the double-value problem. However, this method faces practical limitations, as the size of $\Gamma$ increases with system size, making it unsuitable for larger systems.
Here, we introduce a phaseless SSR(2,2)60, 61, 62 formalism to address both the CI singularities and the double value problem of coupling elements. Our approach is based on deriving the SSR(2,2) equations using the phaseless coupling, $\Delta^{2}$, which represents the square of the off-diagonal elements in the diabatic Hamiltonian, rather than the phase-dependent $\Delta$. We trained equivariant message-passing MLPs to model diabatic PESs and the phaseless $\Delta^{2}$ for the penta-2,4-dieniminium cation (PSB3). This method resolves the phase ambiguity of coupling elements and enhances the accuracy of MLP prediction of NACV in SSR(2,2) ESMD simulations.
2 Method
2.1 Phaseless SSR(2,2) formalism
We employ the SSR(2,2) methodology to calculate excited electronic states. In the SSR(2,2) method, an ensemble of ground and excited microstates provides $(2\times 2)$ diabatic Hamiltonian, incorporating both perfectly spin-paired singlet (PPS) and open-shell singlet (OSS) configurations. These configurations are determined by two Kohn-Sham (KS) orbitals, $\phi_{a}$ and $\phi_{b}$, along with their fractional occupation numbers (FONs), $n_{a}$ and $n_{b}$. From these orbitals, a total of six microstates can be configured, enabling the construction of both PPS and OSS configurations as
$E^{PPS}=\sum\limits_{L=1}^{6}C_{L}^{PPS}E_{L}$ (1)
where $C_{1}^{PPS}=n_{a}/2$, $C_{2}^{PPS}=n_{b}/2$, $-C_{3}^{PPS}=-C_{4}^{PPS}=C_{5}^{PPS}=C_{6}^{PPS}=f(n_{a},n_{b})$ and
$E^{OSS}=\sum\limits_{L=3}^{6}C_{L}^{OSS}E_{L}$ (2)
where $C_{3}^{OSS}=C_{4}^{OSS}=1$, $C_{5}^{OSS}=C_{6}^{OSS}=-1/2$ with KS energy of six microstates, $E_{1}=E^{KS}[\phi_{a}\overline{\phi}_{a}]$, $E_{2}=E^{KS}[\phi_{b}\overline{\phi}_{b}]$, $E_{3}=E^{KS}[\phi_{a}\overline{\phi}_{b}]$, $E_{4}=E^{KS}[\overline{\phi}_{a}\phi_{b}]$, $E_{5}=E^{KS}[\phi_{a}\phi_{b}]$, and $E_{6}=E^{KS}[\overline{\phi}_{a}\overline{\phi}_{b}]$, respectively. $f(n_{a},n_{b})$ is an interpolating function
$f(n_{a},n_{b})=(n_{a}n_{b})^{(1-(1/2)((n_{a}n_{b}+\delta)/(1+\delta)))}$ with $\delta=$0.4. State-averaged energy functional $E^{SA}$ is constructed from $E^{PPS}$ and $E^{OSS}$, by $E^{SA}=w_{PPS}E^{PPS}+w_{OSS}E^{OSS}$. Here, we choose equiensemble condition, $w_{PPS}=w_{PPS}=1/2$. From the optimization of orbitals, we obtain the pseudo-Fock equation,
$n_{p}\hat{F}_{p}\phi_{p}=\sum\limits_{q}\epsilon^{SA}_{pq}\phi_{q}$ (3)
where Fock operator $\hat{F}_{p}$ is given by
$\hat{F}_{p}=\sum\limits_{L}C^{SA}_{L}{\frac{n^{L}_{p\alpha}\hat{F}^{L}_{\alpha}+n^{L}_{p\beta}\hat{F}^{L}_{\beta}}{n_{p}}}$ (4)
Here, $n_{p}$ is an averaged occupation $n_{p}=\sum\limits_{L}C_{L}^{SA}(n^{L}_{p\alpha}+n^{L}_{p\beta})$, $C_{L}^{SA}$ is an averaged weight $C_{L}^{SA}=w_{PPS}C_{L}^{PPS}+w_{OSS}C_{L}^{OSS}$, and $\epsilon^{SA}_{pq}$ is a Lagrange matrix element. Finally, we can construct $(2\times 2)$ secular equation to obtain ground and first excited state energies, $E_{-}^{SSR}$ and $E_{+}^{SSR}$, as
$\begin{pmatrix}E^{PPS}&\Delta^{SA}\\ \Delta^{SA}&E^{OSS}\\ \end{pmatrix}\begin{pmatrix}a_{00}&a_{01}\\ a_{10}&a_{11}\\ \end{pmatrix}=\begin{pmatrix}E_{-}^{SSR}&0\\ 0&E_{+}^{SSR}\\ \end{pmatrix}\begin{pmatrix}a_{00}&a_{01}\\ a_{10}&a_{11}\\ \end{pmatrix}$ (5)
The state interaction between PPS and OSS configuration, $\Delta^{SA}$, is calculated from a Lagrange matrix element and FONs as
$\Delta^{SA}=(\sqrt{n_{a}}-\sqrt{n_{b}})\epsilon^{SA}_{pq}$ (6)
In the SSR(2,2) method, the NACV between two adiabatic states can be directly calculated from the diabatic elements $E^{PPS}$,$E^{OSS}$, and $\Delta^{SA}$ by
$d_{01}=\frac{1}{E_{+}^{SSR}-E_{-}^{SSR}}((a_{00}a_{01}-a_{10}a_{11})\nabla(E^{PPS}-E^{OSS})+(a_{00}a_{11}-a_{01}a_{10})\nabla\Delta^{SA})$ (7)
The coefficients $a_{ij}$ can be expressed with $E^{PPS}$, $E^{OSS}$, and $\Delta^{SA}$ by solving the secular equation. Our key observation is that the NACV can be rewritten based on the phaseless quantity $\Delta^{2}$ as follows:
$d_{01}=\frac{nm}{E_{+}^{SSR}-E_{-}^{SSR}}((-2\Delta^{2})\nabla(E^{PPS}-E^{OSS})+\frac{1}{2}(E^{OSS}-E^{PPS})\nabla\Delta^{2})$ (8)
where $n$ and $m$ represent normalized constants, which have a value of
$n=1/\sqrt{(E_{+}-E^{OSS})^{2}+\Delta^{2}}$ and $m=1/\sqrt{(E_{-}-E^{OSS})^{2}+\Delta^{2}}$, respectively. The significance of Eq. 8 lies in the complete substitution of $\Delta$ with $\Delta^{2}$. By this substitution, we no longer suffer from the phase arbitrariness that is critical for building MLPs for ESMD. We note that the adiabatic state energy and gradient can be rewritten by $\Delta^{2}$ instead of $\Delta$ as
$E_{\pm}^{SSR}=\frac{1}{2}((E^{PPS}+E^{OSS}){\pm}\sqrt{(E^{PPS}-E^{OSS})^{2}+4\Delta^{2}})$ (9)
and
$\nabla E_{+}^{SSR}$ $=$ $m^{2}((E_{-}^{SSR}-E^{OSS})^{2}\nabla E^{PPS}+\Delta^{2}\nabla E^{OSS}+(E_{-}^{SSR}-E^{OSS})\nabla\Delta^{2})$
$\nabla E_{-}^{SSR}$ $=$ $n^{2}((E_{+}^{SSR}-E^{OSS})^{2}\nabla E^{PPS}+\Delta^{2}\nabla E^{OSS}+(E_{+}^{SSR}-E^{OSS})\nabla\Delta^{2})$ (10)
2.2 Machine learning potential for phaseless SSR(2,2)
We employed the equivariant message-passing MLP, NequIP29, implemented in the Bayesian Atoms Modeling (BAM) package, to train three key properties: $E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$. Three independent MLP models were trained to predict each property. The analytical gradients of $\Delta^{2}$, denoted as $\nabla\Delta^{2}$, were computed by expanding $\Delta^{2}$ as $2\Delta\nabla\Delta$. To ensure that $\Delta^{2}$ remained positive, we applied inductive bias layers, such as ReLU activation function, enforcing a non-negative value constraint.
[…]
$\mathcal{L}=\frac{\lambda_{E}}{B}\sum_{b=1}^{B}\left(\hat{E}_{b}-E_{b}\right)^{2}+\frac{\lambda_{F}}{3BN}\sum_{i=1}^{B\cdot N}\sum_{\alpha=1}^{3}\left(-\frac{\partial\hat{E}}{\partial r_{i,\alpha}}-F_{i,\alpha}\right)^{2},$
where $B$ denotes the number of batches, $N$ the number of atoms in the batch, $E_{b}$ the ground-truth energy, $\hat{E}_{b}$ the predicted energy, and $F_{i,\alpha}$ the force on atom $i$ in the direction $\alpha\in\{\hat{x},\hat{y},\hat{z}\}$. The weights $\lambda_{E}$ and $\lambda_{F}$ are set to 1 and 100 (for $\Delta^{2}$, 50), respectively.
The models were trained using the AMSGrad variant65 of the Adam optimizer,66 with the default parameters of $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. We set the initial learning rate to 0.01 and used a batch size of 5. To adjust the learning rate, we applied an on-plateau scheduler based on the validation loss, with a patience of 500 epochs and a decay factor of 0.9. Additionally, we used an exponential moving average with a weight of 0.99 for evaluation on the validation set and for the final model.
Figure 1: Overall schematic for this work. (a) Architecture of the NequIP-BAM model. (b) Global pooling block: ReLU activation enables learning of the phaseless coupling term $\Delta^{2}$(top), while the identity function shows discontinuities (bottom). (c) Nonadiabatic molecular dynamics schematic
The NequIP-BAM model architecture consists of four major blocks: embedding block, interaction block, self-interaction block, and global pooling block. The embedding block maps atomic chemical properties and positional information into vector space, enabling efficient processing by the model. The interaction block models inter-atomic interactions while maintaining rotational equivariance. The self-interaction block updates features by applying identical weights for each atom and uses different weights according to rotational orders to maintain equivariance. Finally, the global pooling block aggregates features from each atom to generate final predictions. In particular, the global pooling block effectively learns the phaseless coupling term $\Delta^{2}$ by applying ReLU activation function, ensuring positive values. The trained model enables nonadiabatic molecular dynamics that can clearly distinguish between $trans$-PSB3 and $cis$-PSB3 isomers, accurately predicting their positions and dynamical behavior on the potential energy surfaces
3 Results and discussion
3.1 Model evaluation
We first evaluated the models for their ability to predict diabatic properties, specifically $E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$. A database was generated consisting of 48,750 data points, randomly selected from 50 surface-hopping trajectories obtained using exact factorization-based dynamics (SHXF)21. These trajectories were computed using the SSR(2,2) formalism combined with the $\omega$PBEh67/6-31g*68 level of theory, with a time step of 0.24 fs over a total simulation duration of 300 fs.
Energy Gradient
MAE RMSE R2 MAE RMSE R2
$E^{PPS}$ 0.041 0.074 0.999 0.065 0.245 0.999
$E^{OSS}$ 0.057 0.102 0.999 0.081 0.314 0.999
$\Delta^{2}$ 0.564 0.978 0.998 1.370 5.840 0.992
Table 1: Mean Absolute Error (MAE), Root-Mean-Square Error (RMSE), and R2 values of the trained NequIP-BAM model. $E^{PPS}$ and $E^{OSS}$ are given in kcal/mol, and $\Delta^{2}$ values are in (kcal/mol)2.
Figure 2: Comparison of diabatic Hamiltonian elements (in eV) between predictions from the NequIP-BAM model and reference SSR(2,2) values: (a) $E^{PPS}$ (b) $E^{OSS}$ (c) $\Delta^{2}$.
As shown in Table 1 and Figure 2, the MLP predictions for diabatic properties are highly accurate. The MAE for the energy predictions of $E^{PPS}$ and $E^{OSS}$ is approximately 0.041 and 0.057 kcal/mol, respectively, while the gradient errors are about 0.065 and 0.081 kcal/mol/Å. Both $E^{PPS}$ and $E^{OSS}$ show high accuracy, surpassing the threshold for chemical accuracy. However, the error for $E^{OSS}$ is slightly larger than that of $E^{PPS}$, likely due to the greater difficulty in mapping excited-state PES from molecular geometry compared to ground-state PES.
Increasing the number of layers in the Interaction Block could potentially enhance accuracy. Indeed, when we tested the model’s accuracy with varying numbers of layers, we observed nearly a twofold improvement in accuracy with the addition of an extra layer for $E^{OSS}$. We also compared the model with different numbers of feature channels, and the results showed that increasing the number of features was less effective compared to increasing the number of layers. However, due to computational cost constraints, we used 5 layers and 64 feature channels in our final model. Supplementary Table S1 presents the results for different layer configurations, while Supplementary Table S2 provides the results for various feature channels. For $\Delta^{2}$, the MAE of the energy is approximately 0.564 (kcal/mol)2 for the NequIP-BAM model. Our approach is notable not only for its accuracy but also for successfully resolving the scattered pattern observed in previous work58, which used a phaseless loss function. This result demonstrates that squaring the double-valued $\Delta$ is an effective strategy for generating a continuous, smooth function that can be efficiently trained. This suggests that the model is no longer affected by the double-valued phase problem and can predict nonadiabatic properties with a similar level of accuracy to conventional models.
Next, we compared the effects of different activation functions applied to the bias layer. During $\Delta^{2}$ training, we observed negative output values from MLP, which are undesirable given the positive nature of the $\Delta^{2}$. Therefore, ReLU and SiLU activation functions were applied during the $\Delta^{2}$ training and these results are presented in Table 2. The ReLU activation function improves the accuracy compared against the model without the activation function. The MAE of $\Delta^{2}$ with ReLU activation is 0.564 (kcal/mol)2, compared to 0.646 (kcal/mol)2 for the model with the identity activation. When the reference value was near zero value, using the identity activation function occasionally resulted in negative predictions for $\Delta^{2}$, which are theoretically impossible. In contrast, the ReLU activation function produced positive predictions due to its bias. This suggests that ReLU is effective in reducing prediction errors for $\Delta^{2}$. While ReLU showed a smaller test error compared to the identity function, the SiLU activation resulted in a significantly larger error of 26.686 (kcal/mol)2. This is likely due to SiLU’s handling of negative values, designed to address the dying ReLU problem, which appears incompatible with the strict positive constraint of $\Delta^{2}$. Based on these findings, we concluded that the ReLU activation function, combined with the NequIP-BAM model, is the most suitable MLP architecture for training PSB3 model. This architecture was applied in all subsequent training and ML-ESMD simulations.
Energy Gradient
MAE RMSE R2 MAE RMSE R2
Identity 0.646 1.239 0.997 1.100 4.785 0.994
ReLU 0.564 0.978 0.998 1.370 5.840 0.992
SiLU 26.686 33.935 0.080 2.568 5.411 0.991
Table 2: Mean Absolute Error (MAE), Root-Mean-Square Error (RMSE), and R2 values of the trained $\Delta^{2}$ model. The units are expressed in (kcal/mol)2.
While overall model performance is important, accurately predicting the minimum-energy conical intersection (MECI) structure is critical due to its central role in nonadiabatic dynamics involving multiple electronic states. The degeneracy of the two adiabatic surfaces at the CI is lifted by two branching plane vectors: the difference gradient vector $\boldsymbol{g}$ and the coupling derivative vector $\boldsymbol{h}$. To assess the model’s accuracy at the CI, we compared the predicted branching plane vectors and MECI structures with their respective references. Figure 3 shows the MECI structures and two branching plane vectors for the CI structure, as predicted by the NequIP-BAM model and calculated by SSR(2,2)/$\omega$PBEh/6-31G*. As shown in Figure 3(a), the MECI structures predicted by the NequIP-BAM model and calculated using SSR(2,2) are nearly indistinguishable. The $\boldsymbol{g}$-vector represents the bond elongation motion of the terminal C=C bond, while the $\boldsymbol{h}$-vector corresponds to the torsional motion of the central C=C bond. These two motions are known to significantly contribute to molecular dynamics at the CI in PSB3, as previously reported69. The predicted vectors at the CI region match the direction and magnitude of the reference values.
Figure 3: MECI structure and its branching plane vectors: (a) Comparison of MECI structures predicted by the NequIP-BAM model (blue) and calculated by SSR(2,2)/$\omega$PBEh/6-31G* (red). (b) difference gradient vector ($\vec{g}$) and (a) coupling derivative vector ($\vec{h}$). The red and blue arrows denote vectors calculated by SSR(2,2)/$\omega$PBEh/6-31G and predicted by the NequIP-BAM model respectively.
Since NACV plays a key role in nonadiabatic dynamics, we evaluated the prediction accuracy of NACV in the vicinity of the CI region by comparing the NequIP-BAM model with the reference SSR(2,2) values. NACVs were calculated using eq 8, with all components computed from the MLP predictions of $E^{PPS}$, $E^{OSS}$, $\Delta^{2}$, and their respective gradients. We divided the NACV test set into two groups: one near the CI region and one away from it (Figure 4). Each group of molecules is composed of 404 and 2,285 molecules, respectively, each containing 42 elements. Thus, Figure 4 includes 16968 elements in the vicinity of the CI and 95970 elements distant from the CI. The division was based on an adiabatic energy difference of 0.5 eV. Both positive and negative vector elements were predicted, matching their reference values, unaffected by the double-valued problem. Accurate NACV prediction near the CI is crucial for describing correct state transitions. Additionally, accurate prediction of NACVs far from the CI region is also important to prevent erroneous transitions in regions where the coupling between electronic states is weak. We confirm that the NACV predictions are not overestimated and remain accurate when it’s away from the CI (Figure 4). In surface hopping dynamics, where hopping probabilities are directly influenced by NACVs, the precise prediction of NACVs by our MLP strengthens the reliability of state transitions in our ML-ESMD simulations.
Figure 4: Scatter plots of NACV predictions from the NequIP-BAM model compared to the reference SSR(2,2) values: Adiabatic energy difference < 0.5 eV (blue) and adiabatic energy differences $\geq$ 0.5 eV (red).
3.2 Nonadiabatic molecular dynamics
After successfully demonstrating that our MLP model accurately reproduces static properties, we ran nonadiabatic dynamics simulations. We used MLP for nonadiabatic MD simulations based on the SHXF method21, implemented in the PyUNIxMD package70, 71. For the initial conditions, 100 nuclear configurations and their velocities were sampled from the optimized ground-state geometry of the $trans$-PSB3 molecule, randomly selected from Wigner distribution on 300K. Dynamics were performed over a total of 300 fs with a timestep of 0.24 fs. The nonadiabatic relaxation of PSB3 from the $S_{1}$ state to the ground state via MECI involves changes in the torsion angle of the central C=C bond, which leads to the differentiation to the $cis$ and $trans$ isomers69.
We tracked the dihedral angle of the central C=C bond (Figure 5). The dihedral angle trajectories clearly show the separation into the $cis$ and $trans$ isomers. Although the results from 100 trajectories may not be statistically significant, the final cis-to-trans ratio of 61:39 is close to the reference ratio of 63:3769. Current work improves on the previous work58, which reported a ratio of 58.5:41.5. Additionally, the comparison of ML-based and SSR-based trajectories reveals that the molecular motions and final yields are in strong agreement between the two approaches (Figure 4 and 5).
We also confirm that the average electronic population evolutions from ML and reference SSR dynamics simulations are in agreement (Figure 5(b)). Averaged populations $P_{i}$ and $p_{i}$ are calculated from the running state and BO population, where $P_{i}=N_{i}/N_{traj}$ and $p_{i}=\sum\limits_{I}^{N_{traj}}\rho_{ii}(t)/N_{traj}$, respectively. Here, $N_{i}$, $N_{traj}$, and $\rho_{ii}$ represent the number of trajectories running on the $i$-th electronic state, the total number of trajectories, and a BO population at time $t$, respectively. The agreement between the two populations highlights the MLP’s ability to accurately capture electron-nuclear correlations. Furthermore, the population trends observed in our SHXF/ML dynamics closely match those from the SHXF/SSR(2,2) simulations, despite the ML model being approximately 100 times faster.
Figure 5: Analysis of the dynamics of photoisomerization of PSB3: (a) Dihedral angle of the central C=C bond of PSB3 over time in individual trajectories. Black trajectories represent ML-based results, while red trajectories represent reference SSR-based dynamics. (b) Average electronic population with the corresponding time. The blue and red line represents ML and reference SSR population evolution, respectively. The solid line represents the BO population and the dashed line represents the averaged running state.
4 Conclusion
We propose a phaseless ML-ESMD approach by building MLPs for the diagonal Hamiltonian elements—$E^{PPS}$, $E^{OSS}$, and $\Delta^{2}$—obtained from the SSR (2,2) method.
This formalism addresses the phase arbitrariness in ML-ESMD. By focusing on predicting diabatic Hamiltonian elements rather than directly on the NACV, we avoid the singularities at CIs. Additionally, we reformulate the NACV and its gradients by replacing the phase-dependent off-diagonal elements with their squared values $\Delta^{2}$. MLP models trained with this approach show high accuracy without any discontinuities or divergence. We show that the ML-ESMD of the PSB3 system agrees well with the reference ab initio simulations. We expect that this approach, when combined with state-of-the-art MLPs, will aid in the development of universal excited-state MLPs. These models could be integrated into de novo molecular design, allowing for the exploration of ever more complex and challenging problems in the field.
5 Code availability
[…]
{acknowledgement}
This research was supported by the National Research Foundation of Korea (NRF) funded by the Korean government (Ministry of Science and ICT(MSIT))(NRF-2023M3K5A1094813, RS-2023-00257666, RS-2024-00455131, NRF-2022R1C1C1010605) and by Institute for Basic Science (IBS-R036-D1). SYW and CWM acknowledge the support from Brain Pool program funded by the Ministry of Science and ICT through the National Research Foundation of Korea (No. RS-2024-00407680). CWM, THP, and SYW are grateful for the computational support from the Korea Institute of Science and Technology Information (KISTI) for the Nurion cluster (KSC-2023-CRE-0502, KSC-2023-CRE-0355, KSC-2023-CRE-0059, KSC-2022-CRE-0424, KSC-2022-CRE-0115, KSC-2022-CRE-0113, KSC-2021-CRE-0542, KSC-2021-CRE-0496, KSC-2022-CRE-0429, KSC-2022-CRE-0469, KSC-2023-CRE-0050, KSC-2023-CRE-0251, KSC-2023-CRE-0261, KSC-2023-CRE-0332, KSC-2023-CRE-0472, KSC-2023-CRE-0501, KSC-2024-CRE-0358). {suppinfo} See the supplementary material for a detailed compilation of the obtained results as well as further data and analysis to support the points made throughout the text.